OpenStudy (anonymous):

Find the area of a rectangular object which has a length of 4√2 inches and a width of 6√6 and explain in complete sentences, how you arrived at the simplified answer and give the final solution in simplified radical form. plz answer my whole question, plz. I'll give a medal.

OpenStudy (anonymous):

Well, the first thing to know is how to calculate the area of a a rectangle. You probably already know this already, but it is length*width. In this case, that means that to get the area, we can simply do \[4\sqrt{2}*6\sqrt{6}\] And from there, you only have to figure out how to simplify it. The simplest thing to do first is to multiply the parts together that match. Notice that 4 and 6 are not under a square root? Well, we can multiply those together normally, getting \[24\sqrt{2}\sqrt{6}\]

OpenStudy (anonymous):

THANK U VERY MUCH!!!

OpenStudy (anonymous):

You're welcome. You can keep going to make it even simpler.

OpenStudy (anonymous):

okay I'll do than, ty again.

OpenStudy (anonymous):

Of course =)